OpenStudy (anonymous):

A 3/8 foot thick board was stacked on top of a 1/3 foot thick board. What was the total thickness of the stack?

OpenStudy (anonymous):

1/8 in 17/24 in 1/6 in 4/11 in answers

OpenStudy (texaschic101):

3/8 + 1/3......what is the common denominator ?

OpenStudy (anonymous):

idk >.<

OpenStudy (anonymous):

im so stupid

OpenStudy (texaschic101):

what does 3 and 8 both go into.....or you could just multiply 8 x 3 to get the answer....so what is it ?

OpenStudy (anonymous):

24

OpenStudy (texaschic101):

do you know how add fractions ?

OpenStudy (texaschic101):

with different denominators ?

OpenStudy (anonymous):

no il try though

OpenStudy (texaschic101):

do you just want the answer ? I amin a good mood

OpenStudy (texaschic101):

or do you want me to teach you ?

OpenStudy (anonymous):

the answer

OpenStudy (texaschic101):

3/8 + 1/3 = 9/24 + 8/24 = 17/24